OpenStudy (anonymous):

How do i evaluate sec(pi/5)? please show steps or give examples so I can work more problems

OpenStudy (anonymous):

you need calculator

OpenStudy (anonymous):

1.23606 pi/2.54 is a close estimate

OpenStudy (anonymous):

the answer is 1.24 but i wasn't sure how i got this answer.

OpenStudy (anonymous):

1.23606 would round up to 1.24 \[\cos \pi/5 = 1/4(1+\sqrt{5})\] I suggest a calculator

OpenStudy (anonymous):

yes i have a ti83plus. thank you